OpenStudy (anonymous):

turn x+y=5 into the formula y=mx+b

OpenStudy (anonymous):

x+y=5---> subtract x to both sides -x -x ---------------- y = -x+5

OpenStudy (nincompoop):

well, how about understanding y=mx+b first we can mechanically convert any equation in all possible forms, but their meaning should be something we have to concentrate on

OpenStudy (anonymous):

ok thankyou